OpenStudy (anonymous):

Find the slope of line that passing through the pair of points (-1,-9)(-3,10)

OpenStudy (nenadmatematika):

formula for the slope of this line is (y2-y1)/(x2-x1)=(10+9)/(-3+1)=19/(-2)=-19/2